OpenStudy (anonymous):

A 5.0 kg bucket of water is raised from a well by a rope. If the upward acceleration of the bucket is 3.0 m/s^2, find the force exerted by the rope on the bucket.

OpenStudy (jamesj):

The bucket has two forces acting on it: gravity and that of the rope. We know that the net force acting on the bucket results in an acceleration of 3 m/s^2. Hence \[ F_{net} = ma = (5.0 \ kg)(3.0 \ m/s^2) = 15 \ N \] Now right now the expression for \( F_{net} \) and solve for the force exerted by the rope.

OpenStudy (jamesj):

@fudgesickle, this works best if you hang around to talk to us, so we can walk you through if you have problems. So let us know if this now makes sense, or where you are.

OpenStudy (anonymous):

You have to subtract the weight of the bucket to find the tension in the rope? T-mg=ma

OpenStudy (anonymous):

T - mg = ma => T = mg + ma => T =5g + 15 So (5g + 15) is the force exerted by the rope on the bucket